OpenStudy (anonymous):

f(x)=e^5x find f'(2).

OpenStudy (ash2326):

We have \[f(x)= e^{5x}\] Let's find f'(x) \[ f'(x)=\frac{d}{dx} e^{5x}\] We get \[f'(x)= 5\times e^{5x}\] We have to find f'(2), Let's put x=2 \[f'(2)=5\times e^{5\times 2}=5\times e^{10}\]

OpenStudy (anonymous):

did you take the derivative of 5x and then apply the power rule to get the 5 *

OpenStudy (ash2326):

Yeah , \[\frac{d}{dx} e^{ax}= e^{ax} \times \frac{d}{dx}(ax)= ae^{ax}\]

OpenStudy (anonymous):

i see. thanks!

OpenStudy (anonymous):

the answer was 110132.33

OpenStudy (anonymous):

e^10 = 22026.46579